OpenStudy (anonymous):

1.6/5 = ?/14 any one now the ? is

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

any one know

OpenStudy (anonymous):

1.6 * 14 = 5x x = 1.6* 14 / 5 = 22.4 / 5

OpenStudy (anonymous):

= 4.48

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

IF YOU ARE GOOD AT ANALYZING POEMS GO TO http://www.twiddla.com/548645

OpenStudy (anonymous):

um.... ya so ? =5

OpenStudy (anonymous):

i analyze poems for a living!

OpenStudy (anonymous):

4.48

OpenStudy (anonymous):

no ? = 4.48

OpenStudy (anonymous):

AMANDA PLEASE GO TO http://www.twiddla.com/548645